Grade 12General Physics

a gun of mass M fires a bullet of mass m with a velocity v relative to the gun. The average force required to bring the gun to rest in 0.5sec is.

laxman
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 

conservation of linear momentum

let the velocity of gun is V1 wrt earth

so velocity of bullet wrt eart is v-V1

m(v-v1) =Mv1

  v1=mv/(m+M)

now distance travel by gun in 0.5 sec

0=v1-a*0.5

a= 2v1     wher a is retardation of gun

so S= v1 *0.5 -1/2 *a *(0.5)2

       =v1 /4

now from work energy theorem total work done is equal to change in kinetic energy

F.S =1/2 M *v12    -0

F.v1 /4 = 1/2 M *v12

F=2Mv1

  =2Mmv/(m+M)

 

Regards

Arun (askIITians forum expert)